OpenStudy (anonymous):

If f(x)=x2+1 and g(x)=3x+1, find 2*f(4).

OpenStudy (anonymous):

A. 18 B. 34 C.65

OpenStudy (babalooo):

What don't you get? and it's x^2 right for f(x). find f(4) then do 2^f(4)

OpenStudy (haseeb96):

just first put 4 at the place of x in the given equation and show me yur work

OpenStudy (haseeb96):

f(4)=(4)^2 +1 =16+1 =17 now 2^f(4)